OpenStudy (anonymous):

Find the other endpoint of a line segment with one endpoint at (-3, 5) and whose midpoint is (5, 2).

OpenStudy (anonymous):

is it 7/2?

OpenStudy (anonymous):

y-coodinate difference= 5-2=3 x-coordinate difference= 3+5=8 thus y-coordinate of endpoint=2-3=-1 x-coordinate of endpoint= 5+8=13 (13,-1)

OpenStudy (anonymous):

that makes more sense! Thank you